CONSIDER THE WAVE FUNCTION Y X , T = 0.015COS 8X-80T CALCULATE THE PARTICLE VELOCITY AND PARTICLE ACCELERATION AT Y 0.01 , 0.4)


1
Expert's answer
2021-09-17T15:32:46-0400

"y(x,t)=0.015\cos(8x-80t),"

"y_0=(0.01,0.4)."


"v_p(x,t)=\frac{\partial y}{\partial x}=\\-8\cdot 0.015\sin(8x-80t),"

"v_p(y_0)=0.058~\frac ms,"


"a_p=\frac{\partial v_p}{\partial x}=\\-64\cdot 0.015\cos(8x-80t),"

"a_p(y_0)=-0.841~\frac m{s^2}."